write an equation for the parabola with the focus (4,0) and vertex (2,0)
Answer by lwsshak3(11628)

write an equation for the parabola with the focus (4,0) and vertex (2,0)
This is a parabola that opens rightwards with equation of the standard form: (y-k)^2=4p(x-h), (h,k)=(x,y) coordinates of the vertex.
..
For given parabola:
axis of symmetry: x-axis or y=0
p=2 (distance from focus to vertex on axis of symmetry)
4p=8
Equation:
y^2=8(x-2)